# Scaling 4D Representations

### v1

@@ Abstract
Scaling has not yet been convincingly demonstrated for pure self-supervised learning from video. However, prior work has focused evaluations on semantic-related tasks – action classification, ImageNet classification, etc. In this paper we focus on evaluating self-supervised learning on non-semantic vision tasks that are more spatial (3D) and temporal (+1D = 4D), such as camera pose estimation, point and object tracking, and depth estimation. We show that by learning from very large video datasets, masked auto-encoding (MAE) with transformer video models actually scales, consistently improving performance on these 4D tasks, as model size increases from 20M all the way to the largest by far reported self-supervised video model – 22B parameters. Rigorous apples-to-apples comparison with many recent image and video models demonstrates the benefits of scaling 4D representations.
@@ 1 Introduction
In this paper, we attempt to answer these questions by revisiting self-supervised learning while looking at performance on five mostly semantics-free tasks: depth estimation, point and object tracking, camera pose estimation, and Something-something v2 [35], an object-agnostic action classification task (this one is more semantic but has been shown to be great at singling out models with good temporal representations). These are tasks that require strong spatial and temporal abilities – they are less about naming things, more about recovering the 3D geometry over time of (4D) scenes. Making an analogy with human vision, we are interested in inspecting dorsal as opposed to ventral stream skills.
Figure 1: We scale self-supervised ViTs on a large video dataset using vanilla masked auto-encoding, then evaluate on tasks requiring accurate 4D representations such as ScanNet depth estimation [17] and Something-something-v2 action classification [35]. Performance keeps improving with model size from 20M to 22B parameters (here with frozen feature evaluation).
We bring together some of the strongest models available – both image (DinoV2 [50], SigLIP [15]) and video models (VideoPrism [80], VideoMAE-v1 [66] and v2 [70], V-JEPA [8]), pretrained with and without language supervision – and evaluate them, carefully ensuring these are apples-to-apples comparisons. We employ for all tasks and models the same attention-based readouts, on top of either frozen backbones or using full finetuning. Our conclusions in all settings are:
@@
    Video Masked Auto-Encoding (MAE) and V-JEPA both perform well. More recent pretraining schemes may result in strong performance on popular classification tasks, such as Kinetics [43], but underperform here.
@@
We opted to dive into MAE and scale up transformer models using it, from 20M all the way to the by far largest self-supervised video model reported so far – 22B parameters, far higher than existing public self-supervised video models (no text) that top out at 1B parameters. We train on a large collection of 170M 30-second videos and, for efficiency, we introduce a scalable, lightweight decoding scheme. We observe consistent overall improvement on the selected tasks as we increase model size, far beyond the 600M parameter setting past which performance saturation has been observed for semantics-centric evaluations [70]. It has been more widely a common belief [^1] Based on personal communication with many elements in the community – unfortunately negative results do not tend to be formally published. in the community that MAE has mediocre scaling properties, to which our work brings counter-evidence.
Contributions. Our contributions are three-fold: a) a re-evaluation of the state-of-the-art from the perspective of 4D scene representation quality; b) three new MAE-VIT models beyond the largest existing 1B ones, having 2B, 4B and 22B parameters, a model family which we term 4DS and plan to publicly release in early 2025; and c) a simple but scalable new decoding scheme using a coarse grid of decoding tokens, enabling us to train the 22B model efficiently.
Paper structure. We go over related work in the next section. We then explain the paper methodology: baseline models, tasks, readouts, evaluation and MAE scaling approach in Sec. 3. Sec. 4 contains all the results and learnings before the paper concludes in Sec. 5.
@@ 2 Related work
Video is a privileged medium for unsupervised learning – a high-bandwidth connection to the world, providing multiple viewpoints of the same objects through time, across lighting changes, deformation and so on. Much early work in self-supervised learning in computer vision focused on video, for example by leveraging temporal continuity signals [32, 76, 48], later boosted by object tracking information [72], motion segmentation [54], camera pose between frames [1], audio [4, 5, 63, 52, 45, 3, 56, 49], future prediction [61, 69, 34], contrastive learning under augmentation [41, 37, 38, 57] and regression-based approaches [62, 30]. Context prediction gave strong results as supervision for text [16], and soon proved effective for visual representation learning [20, 53, 40]. With the rise of Vision Transformers [24], MAE became a particularly efficient formulation of context prediction [39], since masked patches can be skipped by the encoder.
Scaling.
@@
Video versions of masked modeling achieved strong results [66, 8] but the excitement was not long-lasting as attempts to scale models like MAE have yet to succeed [42]. In parallel, CLIP [59] took over, enabling learning from billions of labels from the web, and this translated well also to mainstream video tasks [73, 80] such as classification [43] and action localization [12] (which are defined using words as labels). Models evaluated on the basis of their features (e.g. representations) have been convincingly scaled using language supervision [19] for images. Attempts to scale from just pixels resulted in more modest improvements [13, 26]. In the video space, GPT-like [58] vision and language video models have demonstrated clear benefit from scaling [2]. Pure video self-supervised models have not yet been scaled anywhere near to comparable scales, with some of the biggest models having around 1B parameters [70], or 3B for auto-regressive models [7] – this is something now explored with this paper. Related to our approach, VideoMAE v2 [70] scales a video MAE model, but focuses solely on action recognition benchmarks. They observe only small performance improvements when scaling from ViT-H to ViT-g, and therefore do not scale models beyond the billion parameter regime.
Dorsal streams and evaluation.
@@
The dorsal stream arises in the two-streams hypothesis [47, 33, 31], that the human visual cortex contains two pathways: the ventral stream (which performs object recognition) and the dorsal stream (which is sensitive to transformations and motion). It has been the (indirect) inspiration for several of the early architectures for video representation [64, 29, 28]. Several recent papers studied representations on 3D tasks [25, 79], with results pointing to shortcomings in models solely supervised with language but considered only images. Benchmarks such as Physion [9] and the Perception Test [55] focus on videos and measure more closely the 4D capabilities we are interested in (we adopt one of the tasks from the latter).
@@ 3 Methodology
Our goal is to re-evaluate a large set of strong models apples-to-apples on 4D representation tasks that focus on geometry and motion. To achieve this we follow a representation learning framework where pretrained models get appended with attention-based readouts on top and are then separately finetuned for each of the various downstream datasets / tasks, with and without freezing the backbone. In our experiments detailed in Sec. 4 we found MAE to do well and to be efficiently trainable so we scaled ViT models in that framework – details on scaling are given in the next subsection, followed by the overall evaluation protocol including readout architectures and finally a list of baselines.
3.1 4DS: scaling MAE of video
MAE works by feeding a subset of patches from a video to the video encoder that we want to learn. The features out of the encoder are fed to a decoder that must produce back the missing patches. The decoder is often the expensive part because it needs to produce many more patches (in this paper 95%) than the encoder consumes (in this paper 5%). We employ a simple version of video MAE, hoping that the scale of the models and the data lead to high performance by itself.
We use MAE to train a family of 7 models, that we name 4DS, spanning 20M to 22B parameters and listed in Tab. 1. We use input space-time patch size 2x16x16 for all models.
Data. We pretrain 4DS models using a dataset of 170 million web videos, each 900 frames long on average, at original resolution and frame rate. We sample batches of 16 frame clips from these videos, using temporal stride 2 and randomly cropping down to 224x224 (models up to 4B parameters) or 256x256 resolution (22B parameter models) after resizing the smallest side to 1.15x those resolutions.
@@
Figure 2: Proposed Video MAE framework. Videos are encoded into space-time tokens via one non-overlapping convolution typical of ViTs, then 95% are randomly dropped. A transformer model then processes the remaining tokens with self-attention layers; first alone, then jointly with learned grid of latent tokens on the last few self-attention layers (last two for the 22B model). Each of these learned tokens is decoded back to a patch using a linear layer. We do not use a separate decoder. For downstream evaluations, video features from any layer can be input to additional trainable light-weight readouts.
@@
Decoding. For efficiency and flexibility reasons, we do not use a completely separate decoder with different architectural details, unlike in previous popular MAE approaches [39, 66]; instead the last few self-attention blocks of the encoder share some of the decoding burden but otherwise decoding back to video space is patchwise linear. The process is illustrated in Fig. 2: we concatenate a learned grid of latent tokens to the unmasked set of tokens before one of the last self-attention blocks. We then decode these extra tokens linearly and patchwise, by mapping each to an appropriate number of channels such that they can be reshaped back to the original video dimensions. In downstream tasks, the latent tokens are not used and features can be extracted from any layer. While for models up to 4B parameters we use a decoding grid of tokens that mirrors the original input video grid ($8\times 14\times 14=1568$ (8 for time, 14x14 for space) tokens for 16 frames at 224x224 resolution), for 22B we employ a much more efficient coarser grid ($4\times 8\times 8=256$ tokens for 16 frames at 256x256 resolution). We reconstruct all space-time patches, both masked and unmasked.
@@
(b) Action classification on the SSv2 dataset [35].
(c) Object tracking on the Waymo Open dataset [65].
(d) Camera pose estimation on the RealEstate10k dataset [67].
(e) Depth estimation on the ScanNet dataset [17].
Figure 3: Evaluation downstream tasks. Individual frames and annotations from the five tasks considered in this paper.
Learning. We mask in all cases randomly 95% of the space-time patches (no tricks like tube masking or any special masking pattern), feeding the remaining to the transformer model. The standard AdamW optimizer [46] is used, in conjunction with an L2 loss between RGB values without any special tricks (e.g. no contrast normalization [39, 66]). We train models up to 4B parameters on 2 billion clips, and the 22B models on 4 billion clips, all 16 frames long and sampled randomly from the long 170M videos – smaller models see significantly more data relative to their number of parameters, which benefits them. We hope to train the larger models longer in the future.
Implementation details. We train each model using 256 TPUs-v6, with bfloat16 precision except for the loss and softmax computation which is upcast to float32; weights are stored in float32. We employ simple FSDP-like sharding [51] of model parameters and optimizer state in order to abide by per-device memory limits (32GB). Additional details including all hyperparameters for training are provided in the supplementary material.
@@ 3.2 Evaluation protocol
Tasks. The five tasks we consider are shown in Fig. 3 and capture different aspects of spatial and temporal scene understanding: (a) point tracking on Perception Test [55] measures fine-grained, temporally extended motion perception; (b) action classification in SSv2 [35] is well known to require temporal sensitivity (e.g. shuffling frames leads to very poor performance), (c) object tracking from bounding boxes in the Waymo Open Dataset [65] emphasizes object-level motion. We also include two tasks that connect with 3D perception: (d) 6D camera pose estimation on the Real Estate 10K dataset (RE10k) [67, 81] and (e) monocular depth estimation in ScanNet [17]. All of the tasks are defined in the literature except for the RE10k task for which we introduce a custom metric as well – we give full details about all the tasks including this new one in the supp. material.
Attention-based readouts. There are several common strategies for evaluation of representation learning methods on downstream tasks, particularly with frozen backbones, ranging from non-parametric (nearest-neighbor) [50], linear or using cross-attention layers [8, 80, 66]. In this work we use cross-attention-based readout heads for all tasks: for both SSv2 and RE10k a single learned query is used as in [8], employing an additional step of Procrustes-based projection [11] back into $SO(3)$ for RE10k. For ScanNet we use Fourier-based spatial cross-attention queries, one for each 2x8x8 patch. For the tracking tasks we employ one query per object track for Waymo Open, and one query for each 2-frame temporal chunk of a point track for Perception Test (so 8 queries per track for 16 frame clips). Full details of readout head architectures and hyper-parameters are provided in supp. material.
@@ 3.3 Baselines
Table 2: Comparison of existing (first two blocks of rows) and new models (4DS, last set of rows) when learning an attention readout on top of a frozen backbone. Bold indicates best performance, underlined indicates second-best. All results are obtained using the same evaluation protocol, with exact same inputs, same trainable readouts, and using exact same (limited) number of training examples – see text for details. ScanNet values are multiplied by 10 for better readability.
@@
Model Size (M) Image (I)/ Video (V) Language pre-training SSv2 RE10k Percep. test ($\uparrow$) ScanNet Waymo Open ($\uparrow$)
($\uparrow$) ($\downarrow$) ($\downarrow$)
SigLIP [15] 1,705 I 50.5 2.71 40.2 1.42 45.5
DinoV2-L [50] 303 I 52.2 2.00 36.6 1.03 51.7
DinoV2-g [50] 1,135 I 54.8 1.89 39.9 0.92 50.6
VideoPrism-B [80] 114 V 60.7 0.93 71.9 1.22 66.6
VideoPrism-g [80] 1,113 V 65.4 0.90 75.0 0.98 68.5
V-JEPA-L [8] 307 V 66.0 0.39 75.6 1.14 73.3
V-JEPA-H [8] 635 V 68.9 0.34 78.0 1.05 74.9
VideoMAE-B [66] 87 V 52.3 0.37 79.2 1.50 73.1
VideoMAE-L [66] 305 V 62.7 0.28 78.3 1.10 74.9
VideoMAE-H [66] 633 V 64.2 0.30 77.1 1.04 74.6
VideoMAEv2-g [70] 1,013 V 65.6 0.51 73.7 1.15 72.6
4DS-S 24 V 40.8 0.73 75.9 2.23 69.6
4DS-B 91 V 49.6 0.51 78.9 1.64 72.7
4DS-L 310 V 57.8 0.33 81.5 1.24 75.9
4DS-H 639 V 60.4 0.29 81.8 1.15 76.1
4DS-G 1,848 V 63.1 0.27 82.7 1.02 77.4
4DS-e 3,811 V 64.6 0.27 82.4 0.94 78.0
4DS-j 21,495 V 68.2 0.24 83.4 0.85 78.3
Table 3: Comparison of existing (first two blocks of rows) and new models (4DS, last set of rows) with short finetuning for 20k steps (nothing is frozen). Bold indicates best performance, underlined indicates second-best. All results are obtained using the same evaluation protocol, with exact same inputs, same trainable readouts, and using exact same (limited) number of training examples – see text for details. ScanNet values are multiplied by 10 for better readability.
@@
Model Size (M) Image (I)/ Video (V) Language pre-training SSv2 RE10k Percep. test ($\uparrow$) ScanNet Waymo Open ($\uparrow$)
($\uparrow$) ($\downarrow$) ($\downarrow$)
SigLIP [15] 1,705 I 57.4 2.84 32.9 1.14 44.9
DinoV2-L [50] 303 I 61.0 2.00 48.3 0.91 50.8
DinoV2-g [50] 1,135 I 64.2 2.09 40.5 0.76 50.2
VideoPrism-B [80] 114 V 63.6 0.81 72.8 1.1 66.6
VideoPrism-g [80] 1,113 V 68.6 0.81 73.7 0.89 67.1
V-JEPA-L [8] 307 V 70.4 0.35 76.0 0.96 74.1
V-JEPA-H [8] 635 V 72.9 0.29 71.5 0.87 74.6
VideoMAE-B [66] 87 V 54.8 0.36 79.7 1.51 72.0
VideoMAE-L [66] 305 V 65.6 0.28 79.1 1.05 74.0
VideoMAE-H [66] 633 V 66.5 0.32 77.3 0.93 73.3
VideoMAEv2-g [70] 1,013 V 70.1 0.42 76.7 0.96 71.1
4DS-S 24 V 40.3 0.72 75.2 2.04 69.1
4DS-B 91 V 52.4 0.48 79.5 1.51 71.8
4DS-L 310 V 61.8 0.37 81.1 1.06 74.8
4DS-H 639 V 64.9 0.36 81.9 0.94 75.7
4DS-G 1,848 V 67.4 0.29 83.0 0.81 77.2
4DS-e 3,811 V 68.9 0.28 82.9 0.75 77.5
4DS-j 21,495 V 72.9 0.26 84.2 0.63 78.0
@@
Image models. We evaluate SigLIP-2B which is part of PaLI-3 [15] and a follow-up of SigLIP [78]. This is a large 2B vision transformer model trained from image-text pairs via a binary classification formulation. We also evaluate another extremely powerful and popular image model – Dino-V2 [50] – in two versions, VIT-L and VIT-g, with $303$M and $1.1$B parameters, respectively. Dino-V2 is trained very differently and quite complementary to SigLIP-2B, using a self-distillation loss from images only (no text). SigLIP-2B is one of the strongest descendants of CLIP, Dino-V2 of SimCLR [14]. Both models are trained on large web-scale image datasets.
@@
We also evaluated two types of popular video-only self-supervised models: VideoMAE [66, 70] and V-JEPA [8]. They are both trained on masked videos – VideoMAE is tasked with predicting the pixels, while V-JEPA predicts features encoded by a teacher network. Both models use a ViT backbone and tokenize videos into patches of size (2, 16, 16), applying self-attention between all of the tokens. We evaluate different variants of VideoMAEv1, VideoMAEv2 and V-JEPA, ranging from ViT-B with 30M parameters to ViT-g with 1B parameters (the largest that exist as far as we know).
@@
(a) Reconstructions of the 1st and 9th frames (0.64s apart) from a 16-frame Epic Kitchens [18] video using 4DS models of different sizes (the 22B model is omitted due to differing input/patch dimensions). The last two columns show the original and 95% randomly masked frame prior to reconstruction.
@@
(b) Depth map predictions on ScanNet dataset. White regions in the ground truth indicate missing depth information.

### v2

@@ Abstract
Scaling has not yet been convincingly demonstrated for pure self-supervised learning from video. However, prior work has focused evaluations on semantic-related tasks – action classification, ImageNet classification, etc. In this paper we focus on evaluating self-supervised learning on non-semantic vision tasks that are more spatial (3D) and temporal (+1D = 4D), such as camera pose estimation, point and object tracking, and depth estimation. We show that by learning from very large video datasets, masked auto-encoding (MAE) with transformer video models actually scales, consistently improving performance on these 4D tasks, as model size increases from 20M all the way to the largest by far reported self-supervised video model – 22B parameters. Rigorous apples-to-apples comparison with many recent image and video models demonstrates the benefits of scaling 4D representations. Pretrained models are available at https://github.com/google-deepmind/representations4d.
@@ 1 Introduction
In this paper, we attempt to answer these questions by revisiting self-supervised learning, focusing the evaluation on four semantics-free tasks: depth estimation, point and object tracking, camera pose estimation. These are tasks that require strong spatial and temporal abilities – they are less about naming things, more about recovering the 3D geometry over time of (4D) scenes. Making an analogy with human vision, we are interested in inspecting dorsal as opposed to ventral stream skills. Most prior work has focused on classification downstream tasks such as Something-something v2 [35] and Kinetics [42], so we include results on those as well.
Figure 1: Performance keeps improving on all downstream video tasks as we scale self-supervised video models to up to 22B parameters (shown: frozen feature evaluation). We find that simple scaling of ViTs with masked auto-encoding already delivers for non-semantic video tasks requiring accurate 4D representations such as RE10K camera pose estimation [68] (left, lower is better) and ScanNet depth estimation [17] (middle, lower is better). No prior self-supervised video models have ventured beyond 1B parameters, but they have focused on semantic tasks such as SSv2 action classification [35] (right, higher is better).
We bring together some of the strongest models available – both image (DinoV2 [50], SigLIP [15]) and video models (VideoPrism [82], VideoMAE-v1/v2 [67, 71], V-JEPA [8]), pretrained with and without language supervision – and evaluate them, carefully ensuring these are apples-to-apples comparisons. We employ for all tasks and models the same attention-based readouts, on top of either frozen backbones or using full finetuning.
Our conclusions in all settings are:
@@
    VideoMAE and V-JEPA both perform well. More recent pretraining schemes may result in strong performance on popular classification tasks, such as Kinetics [42], but underperform on the 4D tasks.
@@
We opted to dive into MAE and scale up transformer models using it, from 20M all the way to the by far largest self-supervised video model reported so far – 22B parameters, far higher than existing public self-supervised video models (no text) that top out at 1B parameters. We train on a large collection of 170M 30-second videos using a custom bare-bones MAE approach we call SimpleMAE. We observe consistent overall improvement on all tasks as we increase model size, far beyond the 600M parameter setting past which performance saturation has been observed for semantics-centric evaluations [71]. It has been more widely a common belief [^1] Based on personal communication with many elements in the community – unfortunately negative results do not tend to be formally published. in the community that MAE has mediocre scaling properties, to which our work brings nuance.
Contributions. Our contributions are three-fold: a) a re-evaluation of the state-of-the-art from the perspective of 4D scene representation quality; b) SimpleMAE: a frugal approach to video MAE that removes complexities such as target normalization, tube masking and custom decoders which produced c) three new MAE-VIT models beyond the largest existing 1B ones, having 2B, 4B and 22B parameters, a model family which we term 4DS. Inference code and representative checkpoints are released online at https://github.com/google-deepmind/representations4d, and also include a 90M model distilled from the 4B one.
Paper structure. We go over related work in the next section. We then explain the paper methodology: SimpleMAE, baseline models, tasks, readouts and evaluation in Sec. 3. Sec. 4 contains all the results and learnings before the paper concludes in Sec. 5.
@@ 2 Related work
Origins. Video is a privileged medium for unsupervised learning – a high-bandwidth connection to the world, providing multiple viewpoints of the same objects through time, across lighting changes, deformation and so on. Much early work in self-supervised learning in computer vision focused on video, for example by leveraging temporal continuity signals [32, 78, 47], later boosted by object tracking information [74], motion segmentation [54], camera pose between frames [1], audio [4, 5, 63, 52, 44, 3, 56, 48], future prediction [61, 70, 34], contrastive learning under augmentation [41, 37, 38, 57] and regression-based approaches [62, 30]. Context prediction gave strong results as supervision for text [16], and soon proved effective for visual representation learning [20, 53, 40]. With the rise of Vision Transformers [24], MAE became a particularly efficient formulation of context prediction [39], since masked patches can be skipped by the encoder.
@@
Scaling. Video versions of masked modeling achieved strong results [67, 8] but the excitement was not long-lasting as attempts to scale models like MAE have yet to succeed (there are no published models beyond 1B parameters). In parallel, CLIP [59] took over, enabling learning from billions of labels from the web, and this translated well also to mainstream video tasks [75, 82] such as classification [42] and action localization [12] (which are defined using words as labels). Models evaluated on the basis of their features (e.g. representations) have been convincingly scaled using language supervision [19] for images. Attempts to scale from just pixels resulted in more modest improvements [13, 26]. In the video space, GPT-like [58] vision and language video models have demonstrated clear benefit from scaling [2]. Pure video self-supervised models have not yet been scaled anywhere near to comparable scales, with some of the biggest models having around 1B parameters [71], or 3B for auto-regressive models [7] – this is something now explored with this paper. Related to our approach, VideoMAE v2 [71] scales a video MAE model, but focuses solely on action recognition benchmarks. They observe only small performance improvements when scaling from ViT-H to ViT-g, and therefore do not scale models beyond the billion parameter regime.
@@
Dorsal streams and evaluation. The dorsal stream arises in the two-streams hypothesis [46, 33, 31], that the human visual cortex contains two pathways: the ventral stream (which performs object recognition) and the dorsal stream (which is sensitive to transformations and motion). It has been the (indirect) inspiration for several of the early architectures for video representation [64, 29, 28]. Several recent papers studied representations on 3D tasks [25, 81], with results pointing to shortcomings in models solely supervised with language but considered only images. Benchmarks such as Physion [9] and the Perception Test [55] focus on videos and measure more closely the 4D capabilities we are interested in (we adopt one of the tasks from the latter).
@@ 3 Methodology
Our goal is to re-evaluate a large set of strong models apples-to-apples on 4D representation tasks that focus on geometry and motion. To achieve this we follow a representation learning framework where pretrained models get appended with attention-based readouts on top and are then separately finetuned for each of the various downstream datasets / tasks, with and without freezing of the backbone. In our experiments detailed in Sec. 4 we found MAE to do well and to be efficiently trainable so we scaled ViT models in that framework – details on scaling are given in the next subsection, followed by the overall evaluation protocol including readout architectures and finally a list of baselines.
3.1 Learning from video with SimpleMAE
MAE works by masking out a subset of input patches to an encoder, then having a decoder produce back the masked patches [53]. Previous image MAE implementations additionally employed a) target normalization [39], where target patches are individually normalized; and b) separate lightweight decoders with custom hyperparameters [39]. Video MAE added to the mix c) tube masking [67], where temporal tubes of patches are masked all at once. We do none of these in our training approach – SimpleMAE – isolating instead the impact of the scale of the models and data on performance, without technical bells and whistles. Using SimpleMAE we train a family of 7 models, that we name 4DS, spanning 20M to 22B parameters, listed in Tab. 1.
No special masking. We mask in all cases randomly 95% of 2x16x16 space-time patches (no tricks like tube masking or any special masking pattern), feeding the remaining to the transformer model.
@@ 3.1 Learning from video with SimpleMAE
Figure 2: SimpleMAE framework. Videos are encoded into space-time tokens via one non-overlapping convolution typical of ViTs, then 95% are randomly dropped. A transformer model then processes the remaining tokens with self-attention layers; first alone, then jointly with a learned grid of latent tokens on the last few self-attention layers (last two for the 22B model). Each of these learned tokens is decoded back to a patch using a linear layer. We do not use a custom separate decoder, nor target normalization, nor tube masking. For downstream evaluations, video features from any layer can be input to additional trainable light-weight readouts.
@@
No custom decoder. We use the last few self-attention blocks of the encoder to do decoding work. The process is illustrated in Fig. 2: we concatenate a learned grid of latent tokens to the unmasked set of tokens before one of the last self-attention blocks. We then decode these extra tokens back to pixel space linearly and patchwise. In downstream tasks, the latent tokens are discarded; features can be extracted from any layer. For models up to 4B parameters we use a decoding grid of tokens the same shape as the input video grid ($8\times 14\times 14=1568$, 8 for time, 14x14 for space, for 16 frames at 224x224 resolution). For 22B we use a more efficient, coarser grid ($4\times 8\times 8=256$ tokens for 16 frames at 256x256 resolution). We reconstruct all space-time patches, including unmasked ones.
No target normalization. We use an L2 loss between RGB values without any special tricks (e.g. no contrast normalization [39, 67]).
@@
(b) Object tracking on the Waymo Open dataset [66].
(c) Camera pose estimation on the RealEstate10k dataset [68].
(d) Depth estimation on the ScanNet dataset [17].
(e) Action classification on the SSv2 [35] and Kinetics [42, 65] datasets.
Figure 3: Evaluation downstream tasks. Individual frames and annotations from the five tasks considered in this paper. The majority of previous work has focused on classification, whereas in this paper we do a deep dive into the other four tasks that are less semantic and more about spatial and temporal (4D) scene perception.
3.2 Training setup
We pretrain 4DS models using a dataset of 170 million web videos, each 900 frames long on average, at original resolution and frame rate. We sample batches of 16 frame clips from these videos, using temporal stride 2 and randomly cropping down to 224x224 (models up to 4B parameters) or 256x256 resolution (22B parameter models) after resizing the smallest side to 1.15x those resolutions. Models up to 4B parameters are trained on 1 billion clips, and the 22B models on 2 billion clips, all 16 frames long and sampled randomly from the long 170M videos – smaller models see significantly more data relative to their number of parameters, which benefits them.
We train each model using 256 TPUs-v6, with bfloat16 precision except for the loss and softmax computation which is upcast to float32; weights are stored in float32. The standard AdamW optimizer [45] is used. We employ simple FSDP-like sharding [51] of model parameters and optimizer state in order to abide by per-device memory limits (32GB). Additional details including all hyperparameters for training are provided in the supplementary material.
@@ 3.3 Evaluation protocol
Tasks. We focus on tasks requiring strong 4D representations. They are shown in Fig. 3 and capture different aspects of spatial and temporal scene understanding: (a) point tracking on Perception Test [55] measures fine-grained, temporally extended motion perception; (b) object tracking from bounding boxes in the Waymo Open Dataset [66] emphasizes object-level motion. We also include two tasks that connect with 3D perception: (c) 6D camera pose estimation on the Real Estate 10K dataset (RE10k) [68, 83] and (d) monocular depth estimation in ScanNet [17]. To connect with prior work, we also include (e) action classification in Kinetics-700-2020 [65] and SSv2 [35] – the latter is well known to require temporal sensitivity (e.g. shuffling frames leads to very poor performance). All of the tasks are defined in the literature except for the RE10k task for which we introduce a custom metric as well – we give full details about all the tasks including this new one in the supp. material.
Attention-based readouts. There are several common strategies for evaluation of representation learning methods on downstream tasks, particularly with frozen backbones, ranging from non-parametric (nearest-neighbor) [50], linear or using cross-attention layers [8, 82, 67]. In this work we use cross-attention-based readout heads for all tasks: for SSv2, Kinetics, and RE10k a single learned query is used as in [8], employing an additional step of Procrustes-based projection [11] back into $SO(3)$ for RE10k. For ScanNet we use Fourier-based spatial cross-attention queries, one for each 2x8x8 patch. For the tracking tasks we employ one query per object track for Waymo Open, and one query for each 2-frame temporal chunk of a point track for Perception Test (so 8 queries per track for 16 frame clips). Full details of readout head architectures and hyper-parameters are provided in supp. material.
@@ 3.4 Baselines
Table 2: Comparison of existing (first two blocks of rows) and new models (4DS, last set of rows) when learning an attention readout on top of a frozen backbone. Bold indicates best performance, underlined indicates second-best. All results are obtained using the same evaluation protocol, with exact same inputs, same trainable readouts, and using exact same (limited) number of training examples – see text for details. ScanNet values are multiplied by 10 for better readability. First 4 tasks are the 4D perception tasks we focus on (and where the larger 4DS models do best). We include on the right SSv2 and K700-2020, which are semantic, classification tasks, to connect with prior work that focused on those. 4DS-B-dist-e is a B model distilled from 4DS-e.
@@
Model Size (M) Image (I)/ Video (V) Language pre-training RE10k Percep. test ($\uparrow$) ScanNet Waymo Open ($\uparrow$) SSv2 K700-2020
($\downarrow$) ($\downarrow$) ($\uparrow$) ($\uparrow$)
SigLIP [15] 1,705 I 2.71 40.2 1.47 45.5 50.5 66.4
DinoV2-L [50] 303 I 2.00 36.6 1.02 51.7 52.2 63.5
DinoV2-g [50] 1,135 I 1.89 39.9 0.91 50.6 54.8 65.4
VideoPrism-B [82] 114 V 0.93 71.9 1.20 66.6 60.7 64.3
VideoPrism-g [82] 1,113 V 0.90 75.0 0.96 68.5 65.4 70.3
V-JEPA-L [8] 307 V 0.39 75.6 1.14 73.3 66.0 55.2
V-JEPA-H [8] 635 V 0.34 78.0 1.02 74.9 68.9 57.0
VideoMAE-B [67] 87 V 0.37 79.2 1.51 73.1 52.3 38.9
VideoMAE-L [67] 305 V 0.28 78.3 1.09 74.9 62.7 52.5
VideoMAE-H [67] 633 V 0.30 77.1 1.03 74.6 64.2 54.5
VideoMAEv2-g [71] 1,013 V 0.51 73.7 1.08 72.6 65.6 69.7
4DS-S 24 V 0.73 75.9 2.05 69.6 39.9 28.1
4DS-B 91 V 0.51 78.9 1.65 72.7 49.6 35.7
4DS-B-dist-e 91 V 0.34 81.4 1.21 76.3 60.3 46.4
4DS-L 310 V 0.33 81.5 1.23 75.9 57.6 45.2
4DS-H 639 V 0.29 81.8 1.14 76.1 60.0 47.5
4DS-G 1,848 V 0.27 82.7 1.01 77.4 62.5 52.1
4DS-e 3,811 V 0.27 82.4 0.95 78.0 64.3 54.4
4DS-j 21,495 V 0.24 83.4 0.84 78.3 68.2 58.0
Table 3: Same as Tab. 2 but with short finetuning for 20k steps (nothing is frozen). The trends are consistent with frozen evaluation, but most results are better in absolute terms.
@@
Model Size (M) Image (I)/ Video (V) Language pre-training RE10k Percep. test ($\uparrow$) ScanNet Waymo Open ($\uparrow$) SSv2 K700-2020
($\downarrow$) ($\downarrow$) ($\uparrow$) ($\uparrow$)
SigLIP [15] 1,705 I 2.84 32.9 1.33 44.9 57.4 67.7
DinoV2-L [50] 303 I 2.00 48.3 0.85 50.8 61.0 62.4
DinoV2-g [50] 1,135 I 2.09 40.5 0.72 50.2 64.2 64.8
VideoPrism-B [82] 114 V 0.81 72.8 1.07 66.6 63.6 59.6
VideoPrism-g [82] 1,113 V 0.81 73.7 0.87 67.1 68.6 71.2
V-JEPA-L [8] 307 V 0.35 76.0 0.94 74.1 70.4 56.8
V-JEPA-H [8] 635 V 0.29 71.5 0.85 74.6 72.9 57.6
VideoMAE-B [67] 87 V 0.36 79.7 1.50 72.0 54.8 36.4
VideoMAE-L [67] 305 V 0.28 79.1 0.98 74.0 65.6 50.3
VideoMAE-H [67] 633 V 0.32 77.3 0.93 73.3 66.5 53.0
VideoMAEv2-g [71] 1,013 V 0.42 76.7 0.91 71.1 70.1 70.1
4DS-S 24 V 0.72 75.2 2.05 69.1 40.3 25.4
4DS-B 91 V 0.48 79.5 1.49 71.8 52.4 34.9
4DS-L 310 V 0.37 81.1 1.07 74.8 61.8 45.2
4DS-H 639 V 0.36 81.9 0.96 75.7 64.9 48.6
4DS-G 1,848 V 0.29 83.0 0.81 77.2 67.4 53.0
4DS-e 3,811 V 0.28 82.9 0.73 77.5 68.9 52.7
4DS-j 21,495 V 0.26 84.2 0.62 78.0 72.9 57.3
@@
Image models. We evaluate SigLIP-2B, part of PaLI-3 [15] and a follow-up of SigLIP [80]. It is a large 2B vision transformer model trained from image-text pairs via a binary classification formulation. We also evaluate the extremely powerful and popular image model Dino-V2 [50] in two sizes: VIT-L and VIT-g, with $303$M and $1.1$B parameters, respectively. Dino-V2 is trained very differently and quite complementary to SigLIP-2B, using a self-distillation loss from images only (no text). SigLIP-2B is one of the strongest descendants of CLIP, Dino-V2 of SimCLR [14]. Both models are trained on large web-scale image datasets.
@@
We also evaluated two types of popular video-only self-supervised models: VideoMAE [67, 71] and V-JEPA [8]. They are both trained on masked videos – VideoMAE is tasked with predicting the pixels, while V-JEPA predicts features encoded by a teacher network. Both models use a ViT backbone and tokenize videos into patches of size (2, 16, 16), applying self-attention between all of the tokens. We evaluate different variants of VideoMAEv1, VideoMAEv2 and V-JEPA, ranging from ViT-B with 30M parameters to ViT-g with 1B parameters which are, to the best of our knowledge, the largest that exist.
@@
(a) Reconstructions of the 1st (top) and 9th (bottom) frames (0.64s apart) from a 16-frame Epic Kitchens [18] video using 4DS models of different sizes (the 22B model is omitted due to differing input/patch dimensions). The last two columns show the original and 95% randomly masked frame.
@@
(b) Depth map predictions on the ScanNet dataset. White regions in the ground truth indicate missing depth information.